Four sled dogs are pulling a 1000 kg load. Sled dog A is pulling the load at 20 N, 10° north of east. Sled dog B is pulling the load at 55 N, 70° north of east. Sled dog C is pulling the load at 45 N, 33° north of west, while sled dog D exerts 30 N at a direction of 80° south of west. What is the resultant force acting on the load?

Scale: 1 cm = 10 N




1
Expert's answer
2021-08-31T17:33:00-0400
"F_x=20\cos{10}+55\cos{70}\\-45\cos{35}-33\cos{35}=-4.44\ N\\
F_y=20\sin{10}+55\sin{70}\\+45\sin{33}-30\sin{80}=50.1\ N\\\\F=\sqrt{4.44^2+50.1^2}=50.3\ N\\
\theta=\arctan{\frac{4.44}{50.1}}=5\degree \text{ W of N}"